OpenStudy (anonymous):

solve 9x^4-21x^2+10=0

OpenStudy (saifoo.khan):

This is long one! :(

OpenStudy (jamesj):

Set u = x^2 and then you have a quadratic equation in u, as u^2 = x^4

OpenStudy (anonymous):

factors of 90 that add to -21.....

OpenStudy (jamesj):

are you following what I said?

OpenStudy (anonymous):

discriminant is 81, a perfect square, so it should factor.

OpenStudy (jamesj):

it does factor

OpenStudy (anonymous):

but just use quad formula after a while.

OpenStudy (anonymous):

ok, i got +/-sqrt 5/3 and +/- sqrt 2/3 ???

OpenStudy (anonymous):

\[x= \frac{ 21 +- 9}{18}\]

OpenStudy (jamesj):

(3x^2 - 2)(3x^2 - 5) = 0, yes.

OpenStudy (jamesj):

so pixie yes